# A Quality-Centric Framework for Generic Deepfake Detection

### v2

@@ IV-C Ablation Study
In this section, we conduct ablation studies on different components of the proposed model. All ablation studies are carried out on the HQ of FF++. Note that we aim to verify the effectiveness of the proposed methods targeted on the three research questions. We validate the effectiveness of both the static and dynamic forgery quality assessment strategies for Question 1, as given in Section IV-C4. We validate the effectiveness of the proposed FreDA for Question 2, as given in Section IV-C1. We validate the effectiveness of the proposed curriculum learning framework for Question 3, as given in Section IV-C3.
IV-C1 Scalability and extensibility of FreDA
FreDA augments data through splicing at the frequency domain level, and it can improve any data. Here, we combine FreDA with FF++, Blended images (BIs), and Self-Blended images (SBIs) to improve the data of BIs and SBIs and use these data for model training. The results are shown in Table IV. For the results of BIs and SBIs, we use the code provided by the author of the paper to reproduce. Since their paper uses RAW data for training, this paper uses The HQ version of FF++, so the performance is better than that of the original paper. The results show that the performance of combining BIs and FreDA is 4% higher in terms of AUC than that of BIs on average, the performance of combining SBIs and FreDA is 5% higher than that of SBIs on average, and the performance of combining FF++ and FreDA is 10% higher than that of FF++ on average. And the detection performance of all datasets has been improved. In addition, it is worth noting that BIs and SBIs have fewer forgery traces than FF++, while FreDA has a greater improvement effect on FF++, which also verifies that FreDA has a greater effect on improving samples with obvious forgery traces. This also confirms the rationality of using FreDA on simple samples.
IV-C2 Backbone impact
In Table V, we evaluated the effect of different backbone architectures on the performance of our method, i.e., Xception [12], EfficientNetB4 (EFNB4) [17]. We observe that our method can significantly improve the generalization performances of all evaluated models (at least 9% on average). Meanwhile, as our approach enhances data in a plug-and-play manner, the generalization performance can be improved across different backbones.
TABLE IV: FreDA can improve the generalization of models trained on all types of data.
Methods Celeb-DF Wild DFDC-p DFDC Avg FF++ 74.50 65.45 82.28 68.15 72.60 +FreDA 83.95 76.20 93.95 77.41 82.88 BIs 78.64 71.57 81.10 70.06 75.34 +FreDA 82.28 72.87 91.87 73.99 80.25 SBIs 88.63 63.85 80.01 66.85 74.84 +FreDA 94.79 67.78 89.22 71.27 80.77
TABLE V: Video-level AUC(%) on cross dataset evaluations performances. The best results are highlighted.
Methods CDF Wild DFDC-p DFDC Avg
Xception 69.99 60.06 80.93 65.85 69.21
Ours 82.75 72.29 89.00 70.83 78.72
EFNB4 74.50 61.45 82.28 68.15 71.60
Ours 86.05 77.09 89.95 75.57 82.17
IV-C3 Comparison with other training strategies
In this part, we compared our method with other training strategies, including the vanilla training and BabyStep [62, 57] which is the simplest CL strategy that utilizes a static pre-defined hardness. We conducted BabyStep by introducing the IFQA score as the pre-defined hardness and utilizing the pacing setting in [69].
@@ IV-C3 Comparison with other training strategies
Figure 6: Video-level AUC(%) on cross-datasets performance of different training strategies.
@@
Furthermore, we also investigated the impacts of data augmentation. As shown in Figure 6, we observe that both the CL paradigm and data augmentation can improve performance compared to vanilla training in most cases. However, introducing data augmentation in BabyStep suffers from performance degradation, as it makes the augmented data fail to match its pre-defined static hardness measure. This demonstrates that the dynamic CL strategy with data augmentation (i.e., our method) is beneficial for generic deepfake detection. The key finding here is that the combination of a dynamic CL approach and data augmentation, as used in our method, is more effective than a static CL strategy with data augmentation. This suggests that the flexibility and responsiveness of the dynamic CL approach are crucial for leveraging data augmentation to improve deepfake detection performance.
Figure 7: Visualization of the fake data creation process of FreDA. FreDA enhances the realism (static scores) of the low-quality samples and generates corresponding high-quality samples.
Figure 8: Visualization of BIs and SBIs data processed by FreDA. It can be observed that BIs and SBIs data still have obvious forgery artifacts. After being processed by FreDA, the forged traces in the samples are significantly reduced.
IV-C4 The impact of different definitions of hardness scores
@@
The definition of hardness score plays a crucial role in curriculum learning, so in this section, we use different kinds of hardness scores to conduct ablation experiments. We show the performance of vanilla training (baseline), using only Dynamic Loss , and using all scores (all) as the hardness definition, the results are shown in Table VI. It can be seen that by using DL as the definition of hardness scores, the generalization ability of the model can be greatly improved, achieving a 5.17% gain in terms of AUC than the model obtained by vanilla training. After the introduction of Arcface score, the AUC can be further improved by 3.75%.
After combining all the scores, the performance of the model is further improved, with an average detection performance of 88.49%. It shows that introducing the prior knowledge of and face recognition model (Arcface) helps to improve the definition of hardness score. Arcface score can improve hardness score from different perspectives, which also confirms our motivation to introduce these two static scores.
TABLE VI: The ablation studies on hardness scores.
Methods CDF Wild DFDCp DFDC Avg baseline 74.53 71.56 89.37 71.30 76.69 Dynamic Loss 89.86 72.19 89.85 75.55 81.86 IFQA [16] 93.49 74.29 89.48 75.72 83.25 Ours 94.75 83.77 93.41 82.04 88.49
IV-C5 The impact of the operation on easy samples
@@
Data augmentation is one of the important means of improving the generalization of the model. In this section, we evaluate the impact of different data augmentation on the model’s generalization. We use three settings: no data augmentation (w/o DA), ordinary DA (such as JPEG compression, Gaussian Blur, Affine), and FreDA. The results are shown in Table VII. It can be seen that the generalization ability of the model can be improved by using simple data augmentation, with an average improvement of 2.85%, but there is an obvious decrease in CDF, with a performance reduction of 2.09%. When FreDA is performed on simple samples, the performance is improved across the board, with an average improvement of 7.85% compared to the setting without data augmentation. The results prove that data augmentation has a positive impact on the generalization ability of the model, and also proves the effectiveness of the FreDA we proposed.
TABLE VII: The ablation studies on operation on easy samples.
Methods CDF Wild DFDCp DFDC Avg w/o DA 91.63 70.41 86.37 74.16 80.64 DA 89.54 79.26 89.26 75.89 83.49 FreDA 94.75 83.77 93.41 82.04 88.49
IV-D Visualization
IV-D1 Visualizations of FreDA samples
We randomly screen out four low-quality samples with different ID similarity scores and process them with the FreDA module. The results are shown in Figure 7 and Figure 8. We can see that there are still samples with obvious artifacts of forgery in BIs and SBIs. After being processed by FreDA, the visual quality of FF++, BIs, and SBIs has been improved, and the static score of FF++ has also been greatly improved.
TABLE VIII: Average metrics of different FQS values samples.
Method idx DF F2F FS NT
TAR(%) Top 3.22 3.26 2.32 5.71
Bottom 8.59 9.03 12.32 5.87
SSIM Top 0.9784 0.9932 0.9752 0.9821
Bottom 0.9063 0.9361 0.9376 0.9673
IV-D2 Visualizations of hard samples and easy samples
(a) Real faces with top FQS values
(b) Real faces with bottom FQS values
(c) Fake faces with top FQS values
(d) Fake faces with bottom FQS values
Figure 9: Visualizations of top and bottom FQS values for real and fake faces on FF++(HQ).
We explored the properties of samples with different hardness mined by method. Figure 9 illustrates samples with the highest and lowest FQS values. Fake faces with high FQS values are of relatively high visual quality and difficult to distinguish, while those with low FQS values display clearer forgery clues, such as color inconsistency and blending boundaries. For real faces, samples with high FQS values often undergo heavy post-processing, making them easily mistaken for fake faces. This demonstrates that the FQS values identified by our method align with human visual perception. We also computed the pixel-level tampering ratio (TAR, i.e., the number of tampered pixels divided by the total number of pixels) and SSIM [70] metrics for fake samples with their corresponding real faces. As shown in Table VIII, fake faces with top FQS values involve low TAR and high SSIM, which indicates they have fewer forgery artifacts and higher similarities to the corresponding real faces. It makes sense that deepfake detectors would have difficulty in identifying these samples.
IV-D3 ID similarity difference between FreDA images and original fake images.
Figure 10: The comparison of ID similarity distribution between original fake samples and FreDA samples.
We randomly selected 10 frames from each video, a total of 10,000 images, input these images and their corresponding real images into Arcface, and calculated the ID similarity score. After doing FreDA on these samples, we recalculate the ID similarity score. The comparison of ID similarity distribution is shown in Figure 10. We can see that after processing by the FreDA method, the overall quality of the image has improved, and the mean value has changed from the original 0.560 to the current 0.848, indicating that FreDA can improve the quality of the sample and turn low-quality samples into high-quality samples.
IV-D4 Visualizations of sample hardness changes
Figure 11: Variations of FQS values during the training. We illustrate the change of 3 highest (Data 1-3), 3 lowest (Data 7-9), and random 3 median (Data 4-6) FQS values samples from the 3rd epoch.
Figure 12: t-SNE visualization of latent space with real, original fake (ofake), and FreDA (our augmentation version).
We explored the properties of our training process. We selected samples with the top (hard samples), bottom (easy samples), and median FQS values and illustrated the variations of their FQS values during the training. As shown in Figure 11, we can observe that the FQS values decrease throughout the training for all samples. We also found that the FQS values of easy samples remained small throughout the training. This is because deepfake detectors can learn to identify easy samples in the early stages of training, so our method does not tend to update their FQS. However, the detectors need more time to mine the forged clues of hard samples. This indicates that as learning continues, easy samples become less informative. Selecting and training on fewer hard samples is beneficial, as they provide more valuable information for the model to learn from and improve its performance. Our method allows the model to focus its attention on the hard samples that are more challenging and informative, leading to better overall performance compared to approaches that treat all samples equally or use a static curriculum strategy.
IV-D5 Visualizations of learned latent space.
We input real, original fake (ofake), and FreDA fake samples into a pre-trained deepfake detector, and use t-SNE [71] to visualize the resulting features. The results are shown in Figure 12. It can be seen that the characteristics of fake images processed by the FreDA module are closer to real images than original fake images, indicating that the samples processed by FreDA are of higher quality and closer to real images.
IV-D6 Discussion about Face Reenactment
In this work, the scope for “Deepfake” is specifically the face swapping (FS). However, we also empirically verify that our method can also be applied to face reenactment (FR). Actually, FR is indeed more realistic than FS since FR can be also regarded as two same IDs for face-swapping in its generation process (thus achieving a higher score by our quality evaluator). Results in Tab. II show that our method can also work well for FR.
V Conclusion
@@
This paper proposes a quality-centric framework that explicitly utilizes the forgery quality for generic deepfake detection. The Forgery Quality Score (FQS) is assessed statically via swapping pairs and dynamically based on the model’s feedback. Inspired by curriculum learning, we perform sample selection via FQS during training: a higher quality sample has a higher probability of being sampled for training the model. This manner aims to encourage the model to gradually learn more challenging samples, avoiding overfitting to easy-to-recognizable samples, thereby improving the generalization performance. Furthermore, for these low-quality fake data, we propose a frequency data augmentation method called FreDA, which aims to reduce the easy artifacts of these samples and enhance their realism at the frequency level. Extensive experiments prove that our method can be applied to other approaches in a plug-and-play manner, and can largely improve the generalization of the baseline model. The main limitation of our approach lies in the assumption that obtaining real-fake paired data with the same background is feasible. The assumption is reasonable as in practice, it is possible to conduct such data collection. For example, in the construction of many existing datasets, real and fake images with the same background are often included. This shows that obtaining real-fake paired data with the same background can be achieved in real scenarios. In the future, we aim to explore scenarios where obtaining paired data is not feasible and focus on developing more general forgery quality assessment methods as well as other data augmentation techniques.

### v3

@@ 4.3 Ablation Study
In this section, we conduct ablation studies on different components of the proposed model. All ablation studies are carried out on the HQ of FF++. Note that we aim to verify the effectiveness of the proposed methods targeted on the three research questions. We validate the effectiveness of both the static and dynamic forgery quality assessment strategies for Question 1, as given in Section 4.3.3. We validate the effectiveness of the proposed FreDA for Question 2, as given in Section 4.3.1. We validate the effectiveness of the proposed curriculum learning framework for Question 3, as given in Section 4.3.2.
4.3.1 Scalability and extensibility of FreDA
(a) FreDA can improve the generalization of models trained on all types of data.
@@ 4.3.1 Scalability and extensibility of FreDA
(b) Video-level AUC(%) on cross-datasets performance of different training strategies.
Figure 5: Ablation studies on the proposed FreDA and Pacing Function. (a) Exploration of FreDA’s extensibility; (b) The impact of different training strategies under the cross-dataset setting.
FreDA augments data through splicing at the frequency domain level, and it can improve any data. Here, we combine FreDA with FF++, Blended images (BIs), and Self-Blended images (SBIs) to improve the data of BIs and SBIs and use these data for model training. The results are shown in Fig 5(a). For the results of BIs and SBIs, we use the code provided by the author of the paper to reproduce. Since their paper uses RAW data for training, this paper uses The HQ version of FF++, so the performance is better than that of the original paper. The results show that the performance of combining BIs and FreDA is 4% higher in terms of AUC than that of BIs on average, the performance of combining SBIs and FreDA is 5% higher than that of SBIs on average, and the performance of combining FF++ and FreDA is 10% higher than that of FF++ on average. And the detection performance of all datasets has been improved. In addition, it is worth noting that BIs and SBIs have fewer forgery traces than FF++, while FreDA has a greater improvement effect on FF++, which also verifies that FreDA has a greater effect on improving samples with obvious forgery traces. This also confirms the rationality of using FreDA on simple samples.
4.3.2 Comparison with other training strategies
@@
In this part, we compared our method with other training strategies, including the vanilla training and BabyStep [3, 9] which is the simplest CL strategy that utilizes a static pre-defined hardness. We conducted BabyStep by introducing the static score as the pre-defined hardness and utilizing the pacing setting in [20]. Furthermore, we also investigated the impacts of data augmentation. As shown in Figure 5(b), we observe that both the CL paradigm and data augmentation can improve performance compared to vanilla training in most cases. However, introducing data augmentation in BabyStep suffers from performance degradation, as it makes the augmented data fail to match its pre-defined static hardness measure. This demonstrates that the dynamic CL strategy with data augmentation (i.e., our method) is beneficial for generic deepfake detection. The key finding here is that the combination of a dynamic CL approach and data augmentation, as used in our method, is more effective than a static CL strategy with data augmentation. This suggests that the flexibility and responsiveness of the dynamic CL approach are crucial for leveraging data augmentation to improve detection performance.
4.3.3 The impact of different definitions of hardness scores
@@
The definition of hardness score plays a crucial role in curriculum learning, so in this section, we use different kinds of hardness scores to conduct ablation experiments. We show the performance of vanilla training (baseline), using only Dynamic Quality (DQ), and using DQ in combination with Static Quality (SQ) as the hardness definition. The results are shown in Table 2. It can be seen that by using DQ as the definition of hardness scores, the generalization ability of the model can be improved, achieving a 5.17% gain in terms of AUC than the model obtained by vanilla training. After combining DQ and SQ, the performance of the model is further improved, with an average detection performance of 88.49%. It shows that introducing the prior knowledge of the face recognition model (Arcface) helps to improve the definition of hardness score. Arcface score can improve the hardness score from different perspectives, which also confirms our motivation to introduce static scores.
Table 2: Ablation studies of the proposed modules. (a) Ablations on different hardness scores and (b) ablations of different data augmentations for easy samples with lower hardness scores.
(a) Effect of hardness scores.
Score Test Dataset SQ DQ CDF Wild DFDC-p DFDC Avg 74.53 71.56 89.37 71.30 76.69 89.86 72.19 89.85 75.55 81.86 94.75 83.77 93.41 82.04 88.49
(b) Effect of operations on easy samples.
DA Test Dataset w/o DA Or DA FreDA CDF Wild DFDC-p DFDC Avg 91.63 70.41 86.37 74.16 80.64 89.54 79.26 89.26 75.89 83.49 94.75 83.77 93.41 82.04 88.49
4.3.4 The impact of the operation on easy samples
@@
In this section, we evaluate the impact of different data augmentations on the model’s generalization. We use three settings: no data augmentation (w/o DA), ordinary DA (such as JPEG compression, Gaussian Blur, Affine, or DA), and FreDA. The results are shown in Table 2. It can be seen that the generalization ability of the model can be improved by using simple data augmentation, with an average improvement of 2.85%, but there is an obvious decrease in CDF, with a performance reduction of 2.09%. When FreDA is performed on simple samples, the performance is improved across the board, with an average improvement of 7.85% compared to the setting without data augmentation. The results prove that data augmentation has a positive impact on the generalization ability of the model, and also proves the effectiveness of the FreDA we proposed.
5 Conclusion
@@
This paper proposes a quality-centric framework that explicitly utilizes the forgery quality for generic deepfake detection. The Forgery Quality Score (FQS) is assessed statically via swapping pairs and dynamically based on the model’s feedback. Inspired by curriculum learning, we perform sample selection via FQS during training: a higher quality sample has a higher probability of being sampled for training the model. This manner aims to encourage the model to gradually learn more challenging samples, avoiding overfitting to easy-to-recognizable samples, thereby improving the generalization performance. Furthermore, for these low-quality fake data, we propose a frequency data augmentation method called FreDA, which aims to reduce the easy artifacts of these samples and enhance their realism at the frequency level. Extensive experiments prove that our method can be applied to other methods in a plug-and-play manner, and can largely improve the generalization of the baseline model.